A 228V constant potential generator delivers load to a variable impedance whose values of R and X 

are readily adjustable. Assuming a looking-back impedance of (6+j4.5) ohms calculate, for maximum 

power,

a. The load impedance

b. The current

c. Load power

d. power factor


1
Expert's answer
2021-05-31T06:21:55-0400

a) For maximum power transfer the load impedance "=6-j4.5"


b) The current "=228\/12= 19A"


c) The load power"=I^2R=19^2\times6=2166W"


d) The load power factor is "Cos(tan^{-1}(4.5/6) )=cos36.86\degree=0.8lead"